OpenStudy (anonymous):

How to show your work of completing the square of x^2-12x+7=0

OpenStudy (mathmale):

Example: complete the square of x^2 + 2x + 5. 1) Take half of the coefficient (2) of x. This is 2/2 = 1. 2) Square that quantity 1^2+1. 3) add the result to x^2 + 2x, and then subtract the same thing, and then add that 5: x^2 + 2x +1 -1 + 5 = x^2 + 2x + 1 +4 4) You'll have a perfect square on the left. Write it as the square of a binomial: (x+1)^2 + 4. That's it. Now please apply this procedure to your homework problem.